OpenStudy (anonymous):

Arrange the following polynomial into descending order for x, then interpret the degree of the 2nd term. 7x3y3 + 4 − 11x5y2 − 3x2y

OpenStudy (anonymous):

@TuringTest @mlb004 @ash2326 @whpalmer4 @JANELLSM98

OpenStudy (anonymous):

@jigglypuff314

OpenStudy (anonymous):

@eliassaab

OpenStudy (anonymous):

I know its not 5

jigglypuff314 (jigglypuff314):

what order would the x values be in from greatest to least x^3 x^5 x^2

OpenStudy (anonymous):

x^5 x^3 x^2

jigglypuff314 (jigglypuff314):

yes, so that's how you would rearrange the overall problem

OpenStudy (anonymous):

so from these choices: A. 7 B. 6 C. 5 D. 3 Would it be d? @jigglypuff314

jigglypuff314 (jigglypuff314):

I think so :) but I'm not entirely sure :P you might want to check that part with someone else

OpenStudy (anonymous):

@lovable34 @sourwing

OpenStudy (anonymous):

@BTaylor

OpenStudy (anonymous):

@Loser66

OpenStudy (gorv):

descending order of x means arrange x from highest power to lowest -11x5y2 +7x3y3 − 3x2y +4

OpenStudy (gorv):

degree of second term is 6

OpenStudy (btaylor):

First, we must arrange in descending order: \[7x^3y^3+4-11x^5y^2-3x^2y \rightarrow -11x^5y^2+\color{red}{7x^3y^3}-3x^2y +4\] The degree of that term is the sum of the degree of x and y, so 3+3=6

OpenStudy (loser66):

I don't understand the question, the "interpret the degree of the second term

OpenStudy (gorv):

as powerof x+power of y=degree degree=3+3=6

OpenStudy (anonymous):

@gorv @BTaylor ok is that it?

OpenStudy (gorv):

yeah that's it

OpenStudy (anonymous):

@gorv thank you